Given that 0 (theta, sorry don't have this symbol on my keyboard) is an acute angle with
cos 0 (theta)=19/45'
Find the exact value of cosec 0 (theta)
Thank you for your help

Answer by stanbon(75887)

Given that 0 (theta, sorry don't have this symbol on my keyboard) is an acute angle with
cos 0 (theta)=19/45
Find the exact value of cosec 0 (theta)
-----
cos = x/r, so x = 19 and r = 45
----
The y = sqrt[45^2-19^2] = 40.79
-----
Ans: csc(t) = r/y = 45/40.79 = 1.10
-----------------------
Cheers,
Stan H.
--------------